The length of a rectangle is 10 less than 3 times its width . The perimeter of the rectangle is at most 180 inches. Find the maximum length of the rectangle.
Answer by addingup(3677)

L= 3W-10
perimeter of the rectangle:
2L+2W Substitute in this formula the value for L given in the problem:
2(3W-10)+2W ≥ 180
6W-20+2W ≥ 180 Add W on left and add 20 on both sides:
8W ≥ 200
W ≥ 25 is the width. And the length is:
3(25)-10= 75-10= 65 is the length